QUESTION-2 Consider another Ricardian example, i.e., Comparative Advantage. What David Ricardo saw was that it could still be mutually beneficial for both countries to specialize and trade. Please answer the below given questions for the below given data set for Moonited Republic and Vintland Republic. SHIRTS Cost Per unit man hours 7.5 SHOES Cost Per unit man hours Moonited Republic (6000 man hours) Vintland Republic (2000 man hours) 2.5 (A) Which country has an absolute advantoge in shirts? Explain with numbers. (B) Which country has an gbsolute advantoge in shoes? Explain with numbers (F) Show the NO-TRADE production point for each country SHOES Cost Per unit man hours SHIRTS Cost Per unit man hours 3,000 3,000 Moonited Republic (6000 man hours) Vintland Republic (2000 man hours) 1,000 1,000 (C) Which country has a comparative advantage in shirts? Explain with numbers. (D) Which country has comparative advantage in shoes? Explain with numbers SHIRTS / SHOES | Δ SHOES / Δ SHIRTS (E) Produce the Production Possibilities Curves for both countries. Determine which country has a comparative advantage in shirt? Which country has a comparative advantage in shoes? (G) If trade is opened, which country should produces which good? (H) Show the FREE-TRADE production point for each country. Compare the result with NO-TRADE option with SPECIALIZATIOIN & FREE-TRADE case SHIRTS Cost Per unit man hours SHOES Cost Per unit man hours Moonited Republic (6000 man hours) Vintland Republic (2000 man hours) TOTAL
Answer #1

Monotonic republic can produce either (6000/7.5) = 800 shirts or (6000/3) = 2000 shoes.

Vintland republic can either produce (2000/2.5) = 800 shirts or (2000/2) = 1000 shoes.

A) So with these figures , we can say that none of the countries have an absolute advantage in the production of shirt over the other as both can produce a maximum of 800 shirts individually.

B) With the figures given, we can say that Monotonic Republic has an absolute advantage over the Vintland Republic in the production of shoes as Monotonic Republic can produce a maximum of 2000 shoes with its resources whereas Vintland Republic can only produce 1000 shoes with its resources.

C) Opportunity Cost of producing shirts in Monotonic Republic = 2000/800 = 2.5 shoes.

Opportunity Cost of producing shoes in Monotonic Republic = 800/2000= 0.4 shirts.

Opportunity Cost of producing shirts in Vintland Republic = 1000/800 = 1.25 shoes.

Opportunity Cost of producing shoes in Vintland Republic = 800/1000 = 0.8 shirts.

So we can say that Vintland Republic has a comparative advantage in the production of shirts as the opportunity cost of producing shirt in Vintland Republic is 1.25 shoes which is less than the opportunity cost of producing shirt in Monotonic Republic ,which is 2.5 shoes.

D) With the calculations made above, we can say that Monotonic Republic has a comparative advantage in the production of shoes as the opportunity cost of producing shoes in Vintland Republic is 0.8 shirts which is more than the opportunity cost of producing shoes in Monotonic Republic , which is 0.4 shirts.